Meeting Information

  • 1600 UT (1000 MDT)
  • AOC 317; IP: 146.88.16.8; phone: +1 575-835-7370

Agenda/Minutes

News, Travel, Visitors (McMullin)

  • Overall status: C configuration ready and observing (pointing and collimation are complete; baselines not yet fully determined). Restrictions on time averaging have not yet been relaxed (4 seconds only). SKA meeting in ABQ through today; on a tour of the EVLA on Wednesday. PASEO response due this week; OSRO 2011 D configuration capabilities finalized this week.
    • Report on ALMA Commissioning status (McMullin)
  • Key dates:
  • Staff Travel:
    • Sowinski: 20 Sep - 15 Oct
    • Perley: 23 Sep - 19 Oct
    • van Moorsel: 11-15 Oct
    • Wrobel: 24 Oct-27 Oct
    • DRAO: Dougherty: 11 Oct - 12 Oct (SKA; ABQ)
  • Visitors:
    • Brent Carlson: 18 October - 25 October; second trip post SKA meeting TBD
    • DRAO: Dougherty: 12 Oct - 17 Oct (NRAO-DSOC)
    • Adam Deller: 09 Oct - 16 Oct (phased array testing)
    • Joe Lazio: 12-15Oct (phased array testing)

Hardware status and updates (Langley, Perley, Rupen)

  • Array (Perley/Langley):
  • WIDAR (Rupen):
    • Notes: 3-bit tests on sky have revealed inconsistencies with the 8-bit and with switched power measurements (y-factor tests at different stages along the signal path). More testing this week but proposed actions deal with hardware changes at different levels of severity and time commitment; more by the end of the week. Additional tests looking at the T304 power (3-bit vs 8-bit; sampler and correlated power) show a departure from the expected behavior for 3-bit that is being pursued. Current CBE restriction on time averaging remains at 4s; some OSRO projects let through without modifying (will at least lose the last scan); Martin is working on this Monday-Tuesday to relax the constraint; Update: New version went in at noon; initial tests look good; detailed tests took place last night. CM testing to enable recirculation (desired by some RSRO projects). CRM work ongoing but still not in place for tracking the hardware components and re-routing around failures as needed. Mark 5c testing proceeding this week; dedicated phased array testing Thu-Fri - Update from Adam here: We are recording good packets from WIDAR onto the mk5c via the 10 GbE connection. The initial round of data last week looked bad (bad headers) because the value used for packet offset (to strip the UDP headers) was apparently incorrect - we need to use 32, not 40. Getting some bad packets between good which seems to be coming from the switch; trying to ignore those. More testing today. Update to the RXP binary for phased array went in also.
      • Update on time averaging: tests at 10,30,60 s taken last night; there are issues with averaging for reference pointing (given the 10s subscans used for the pointing positions and the heuristic within TelCal which rejects adjacent scans if one is flagged, it is very easy to lost a chunk of the pointing and defeat the fit); please keep it to 1 s integrations.
  • Computing (Robnett):

Software status and updates (Butler)

  • Links:
  • Notes: New version of the Executor (mostly cleanup+work for phasing the array); MCAF schema changes appear to be handled well; moving forward with version (possibly Wednesday); OPT 1.06 under webtest; archive now working for 10B projects(!).
  • CASA Update (new functionality available/issues/enhancements):Release due Friday; next development cycle planning progressing, to be discussed in a broader forum next week possibly.
  • AIPS Update (new functionality available/issues/enhancements): Work on Bill's code (CL table issues noted by Laura); had trouble with data labeled as lower SB, now using the frequency increment and not trusting the label; continuing to iron out issues in the SDM. Noted that CASA has some information that is not in the SDM (polar coordinates); not clear how these are determined (need answer from Jeff); they don't look like they are being employed; ultimately these will be in the SDM. A range of new tasks to support switched power analysis - EDITA, SNEDIT handle the SY table; need a version of TYAPL for SY tables; TYSMO allows clipping/smoothing of SY tables.

Observing recommendations update (Ops)

Commissioning priorities for the week (McMullin, Chandler)

Somewhat complex this week as there is more of a day-to-day breakdown due to the dedicated campaign Thu-Fri:
  • Mon-Tue: CBE/switched power/CM testing (in that order of priority).
  • Wed: Maintenance day
  • Thu-Fri: Phased array testing

  • Interrupt status for additional 3-bit testing

Commissioning Reports

Array: Dhawan: C configuration move status

  • Move completed early as expected. Pointing updated, baselines updated for all antennas; collimation run will be done when EA14 is ready. Further Rx checkout this afternoon. Seeing some signs of phase jumps in EA16 (on one IF). Have been doing science observing as selected by the OST; RSRO not selected as high frequency wasn't ready until the updates to pointing, etc had gone in; took first RSRO last night (AL746).

Correlator: McMullin/Ott: MCAF update (compare test and production)

  • Tests run on the parallel production and test version, comparing the visibilities through to imaging; looks good so will be updated shortly. Note, that Bill likely hasn't had communication on this update so action for Bryan to communicate this to him.

Planetary: Butler: Venus/TNO observations

  • See https://bugs.nrao.edu/browse/ECSV-109; noted that some problems with tacking (delay and delay rate even for known ephem objects); plan was to point at a fixed position and fix the phase center in post-processing using AIPS (common case used for comets whose orbits are not yet well determined). Observation of Neptune and Triton, pointing at Triton. Doing a normal calibration and averaging; key issue is the time variability imposed by the motion of Neptune. Bottom line is that the technique works; found higher noise but explained in the calibration process which can be improved (slicing in time). Next time will also point at Neptune to avoid the issues which dominated in the first test. Pluto will be tried next.

In preparation

  • Polarization: Multi-band leakage/PA (Myers, Marvil): Not a priority but still needed; keep on list.
  • Correlator: Krauss/Hunter: 7 vs 8 node observing (follow-up; pending RSRO results)

Commissioning Test Planning

  • OPT: webtest of 1.06 capabilities/bug fixes (*); ongoing (Lorant, Michael, Joe)
  • Calibration: Switched power on/off 1dB attenuators. (planned for this week).
  • Array: RF switch tests with rapid monitoring (as possible but likely not this week).
  • Correlator: CBE averaging with restricted nodes
  • Correlator: Quadrant 2 tests; CM update test.
  • Calibration: Switched power tracking over several airmasses.
  • Polarization: Antenna cross polarization (Perley; Need script.); LSCX on 3C147; Delayed until Rick's return
  • Spectral Line: Dhawan: Follow-up on frequency offset phase jumps characterization.
  • Spectral Line: Ott: Doppler tracking sampling follow-up
  • Spectral Line: Momjian: Zeeman test follow-up - inconsistencies have complicated the conclusions. Likely need a small forum to discuss within the spectral line group (and others interested); the Thursday meeting spot is available as needed.

RSRO/ECSO Reports

10B-124: Kellerman: Test run status; being looked at.

10C-219: Irwin: Test run?

AC982: Claire noted some impressive RFI in K band that wasn't originally understood. In addition, plots illustrating delay clunking were shown. See RSRO-4.

AT380: Satoko noted some issues with mis-labeling of scans in her tests which should be handled by some script processing.

Other updates:

  • Notes:
    • Current Residents
      • Volker Heesen (POC: Michael Rupen); area(s): polarimetry
      • Peter Hofner (POC: Juergen Ott); area(s): post-processing
      • Russ Taylor (POC: Rick Perley); area(s): polarimetry
      • Satoko Takahashi (POC: Juergen Ott); area(s): continuum, mosaicing; 10B-126 (AT380): A Millimeter Study of the Embedded Star-Cluster in the Orion Molecular Cloud.
        • Hauyu Liu (POC: Juergen Ott); area(s): continuum, mosaicing
      • Greg Hallinan (POC: Kumar Golap); area(s):; 10B-209 (AH1020): Broadband Periodic Dynamic Spectra of Ultracool Dwarf Pulsars
      • Adam Leroy (POC: TBD); area(s):; 10A-196 (AL746): Resolving the Starbursts in Nearby LIRGs and ULIRGs
      • Jill Knapp (POC: Rupen/Claussen); area(s): recirculation; IRC+10216 demo science; 10B-200;10B-186. Welcome again.

Demo Science (Chandler/Condon): Project Codes: *TDEM*

Project Status Time Comments SB
Crab nebula Complete 10/10 L-band OSRO1/OSRO2 Complete
IRC+10216 Ready (HF) 21/33 Partial; revisit in DnC/C Partial Additional data taken!
M87 Ready (HF) 0/6 Ka Not done
M51 Ready (MR) 8/8   Complete
Cygnus A Ready 3x6 1x10 possible; 12 OSRO hours obtained Withdrawn
M82 Complete 2/2 L band Complete
Herc A Ready 3/5 CX bands Partial

Commissioning Group updates

AOB

Misc Actions

  • [McMullin/Frail]: %REDWrite up disposition of RR items.
  • [McMullin]: Confirm restriction of RSRO scan length at 90 s
  • [Owen]: Outline issue with 90s scan length restriction.
  • [McMullin/Durand]: Clean up outdated MainSaver issues.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/McMullin]: SPW/IF/BB lexicon and mapper
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.
-- JosephMcMullin - 11 Oct 2010
Topic revision: r2 - 2010-10-13, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback